0. given f(x) =8x -2 and g(x) = 2x -3 find f-g
1. given f(x) =5x + 4 and g (x) =3x-8 find f.g


2. given f(x) +2 -2x and g(x) = -6 +2, find f+g


3. find the distance between the pair of points (-1, -4) and (-5, -7)


4. find the distance between the pair of points (-2, -6) and (7, -3)


5. find the midpoint of the line segment with the given and points (2,1) and (6,8)


6. find the center and radius of a circle given the following equation of the circle x^2 + y^2 = 100


7. given f(x) = -3x + 2 and g(x) = 2x +9, find (gof)(x)


8. given f(x) = 5x +7 and g(x) = 5x -1 find (fog)(x)


9. write the standard form of the equation of the circle with the given center(-2, -4) and the radius 6.

Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(24713)



Sorry. We're not going to do your entire homework assignment for you.


John


Answer by Alan3354(69443)

0. given f(x) =8x -2 and g(x) = 2x -3 find f-g
1. given f(x) =5x + 4 and g (x) =3x-8 find f.g

2. given f(x) +2 -2x and g(x) = -6 +2, find f+g

3. find the distance between the pair of points (-1,-4) and (-5,-7)
s^2 = diffy^2 + diffx^2
s^2 = (-4+7)^2 = (-1+5)^2
s^2 = 9 + 16 = 25
s = 5
--------
4. find the distance between the pair of points (-2,-6) and (7, -3)
Do it like #3 above.
----------------------
5. find the midpoint of the line segment with the given and points (2,1) and (6,8)
Get the averages of x any y separately.
for x: (2+6)/2 = 4
for y: (1+8)/2 = 9/2
Midpoint is (4,9/2)
---------------
6. find the center and radius of a circle given the following equation of the circle x^2 + y^2 = 100
A circle of radius r about center (h,k) is:
(x-h)^2 + (y-k)^2 = r^2
This one is centered about the Origin (0,0) with a radius of 10 (sqrt(100)).
-----------------
7. given f(x) = -3x + 2 and g(x) = 2x +9, find (gof)(x)

8. given f(x) = 5x +7 and g(x) = 5x -1 find (fog)(x)

9. write the standard form of the equation of the circle with the given center(-2, -4) and the radius 6.
See #6